# QNM families: classification and competition

### v1

@@ Abstract
The perturbation spectra of black hole (BH) beyond standard general relativity (GR) may exhibit complex structures with long-lived modes that generate echo-like modulations on the ringdown signal. However, a systematic framework for understanding the internal structure of such spectra, the physical nature of different mode families, and their dynamical excitation is still undeveloped. In this paper, we address this issue by proposing a potential methodology that combines frequency-domain classification with time-domain analysis, using a hairy Schwarzschild BH that admits a double-peak perturbative potential as a theoretical platform. Our analysis of the quasinormal mode (QNM) spectrum identifies two distinct families of modes: the photon sphere (PS) family, arising from delocalized scattering resonances, and the echo family, corresponding to highly localized quasi-bound states. We then develop a windowed energy analysis framework in the time domain, which discloses a dynamic competition for dominance between these families. In particular, our results explicitly show that this competition is sensitive to the properties of the initial perturbation source, and that higher-overtone echo modes can dominate in the observed signal, which are in contrast to the standard PS mode in GR. This study establishes the dynamic evolution of this energy competition as a new observational signature for probing new physics and further motivate a supplemental framework for analyzing long-lived ringdown signals.
@@ I Introduction
The dawn of the gravitational-wave (GW) astronomy era, inaugurated by the landmark observations of the LIGO collaborations LIGOScientific:2016aoc , has opened an unprecedented window into the strong field regime of gravity. Future space-based detectors such as LISA Finn:2000sy , Taiji Hu:2017mde , and TianQin TianQin:2015yph and more promise to expand this window even further. One important scenario in this field is black hole spectroscopy Berti:2009kk ; Konoplya:2011qq ; Berti:2025hly , the idea that by analyzing the “ringdown” signal of a post-merger BH, one can probe its intrinsic properties analogous to how atomic spectroscopy reveals the structure of atoms. Within the standard paradigm, the ringdown waveform is modeled as a superposition of QNMs. These modes, characterized by their complex frequencies, are considered carriers of information about the final BH’s mass and spin, providing a powerful framework to test the foundational no-hair theorem of GR Berti:2005ys ; Dreyer:2003bv . This standard picture, however, becomes significantly more complex when considering that the final compact object may not be a Kerr BH, and its perturbation spectrum could exhibit a much richer structure.
@@
Indeed, recent theoretical explorations have shown that the emergence of new QNM families is a widespread phenomenon in a variety of physical scenarios. These new families can be categorized by their physical origins. Firstly, modifications to spacetime’s asymptotic or topological structure can introduce new QNM families. A prime example is found in Cardoso:2017soq , where, in addition to the familiar photon sphere (PS) family tied to the BH’s light ring, a distinct and long-lived de Sitter family arises from the presence of the cosmological horizon. Exotic compact objects like wormholes Cardoso:2016rao ; Bueno:2017hyj ; Qian:2025lau , or even specific braneworld models Tan:2024qij , can feature effective potential wells that trap perturbations, giving rise to long-lived echo families Cardoso:2019rvt . Additionally, the echo modes were also found and analyzed by imposing a non-smooth effective potential Liu:2021aqh ; Daghigh:2025wcw ; Xie:2025jbr . Secondly, the coupling between gravitational perturbations and other matter fields invariably leads to mixed spectra. For examples, the gravito-electromagnetic coupling in Kerr-Newman spacetimes splits the QNM spectra into PS and near-horizon families Dias:2015wqa ; and the gravitational and scalar modes in various scalar-tensor theories become intertwined Guo:2022rms ; Blazquez-Salcedo:2016enn ; Lin:2025lsz . Lastly, the intrinsic properties of the perturbing fields themselves, such as the mass of a Proca field, can generate quasi-bound states that form their own spectral families Rosa:2011my ; Lei:2023wlt . All these studies suggest that complex spacetime geometries or matter contents could leave their imprints directly on the global structure of the QNM spectra.
@@
Our methodology consists of a two-step workflow. The first step is a physical classification of modes in the frequency domain, which is made by WKB approximation, and we verify the reliability of the results of the QNM spectra with a set of methods. The second step is a dynamical analysis in time domain. To this end, we compute the complex amplitude of the master scalar Berti:2007dg ; Nollert:1998ys and develop a windowed energy analysis framework. By calculating the energy fraction of each mode as a function of parameters related with a sliding observation window, we construct a dynamic portrait of how different families compete for dominance and reveal the sensitivity of this competition to the initial perturbation source.
@@ II Perturbation equation and echo parameter space
$\text{with}\quad A(r)=B(r)=1-\frac{2M}{r}+\alpha\,e^{-\frac{r}{M-l_{0}/2}}\equiv f(r)\quad\text{and}\quad C(r)=r^{2}\,$, (2)
This black hole was found to admit an additional PS in certain parameter regimes Guo:2022ghl , comparing against the single PS in Schwarzschild black hole. These multiple PSs bring in richer and distinctive optical appearances Meng:2023htc and interferometric signatures Wang:2025hzu . The presence of additional PS has also found to closely associated with the echo signals of the probe scalar field Cavalcanti:2022cga ; Yang:2024rms . Moreover, the QNM and greybody factor for various perturbations of this black hole has been studied in Yang:2022ifo , in which the authors focus on the case with single peak in the perturbative potentials.
@@
where the field strength tensor $F^{\mu\nu}$ is defined in terms of the vector potential $A_{\mu}$ as $F=\mathrm{d}A$. In a spherically symmetric background, the vector potential can be decomposed into a Regge-Wheeler-Zerilli(RWZ) form
@@
For polar modes with even parity $(-1)^{\ell}$, the master scalar can be taken as
@@
We move on to investigate the morphology of the effective potentials. A crucial aspect of the hairy black hole solution (1) is the rich phenomenology that arises from its full parameter space. While the analysis in Cavalcanti:2022cga ; Avalos:2023jeh focused on a specific subspace by imposing the constraint $l_{0}=2M\alpha/(e^{2}+\alpha)$ which adjusts the horizon radius of the hairy black hole horizon radius to that of Schwarzschild black hole (e.g. $r_{h}(M,\alpha,l_{0})=2\mathsf{M}_{s}$). Here we shall explore the broader implications by treating the hairy charge $l_{0}$ and the deformation intensity $\alpha$} as independent variables. The interplay between these two parameters can induce a non-trivial structure in the spacetime geometry. Specifically, for certain regions in the $\{\alpha,l_{0}\}$ parameter space, the metric function $f(r)$ develops two inflection points outside the horizon. This non-monotonicity of the gravitational potential is directly imprinted onto the perturbative dynamics, causing the effective potential $V_{s}(r)$ to exhibit three local extrema. We then scan the parameter $\{\alpha,l_{0}\}$ and map the morphology of the electromagnetic potential with fixed $s=1$, $\ell=6$ and $M=1$ (same settings for later studies) in Fig. 1, where the grey shaded region highlights the parameter combinations that give rise to the double-peak potential, in contrast to the blank region corresponding to the typical single-peak case.
@@
Figure 1: (a): Parameter spaces of the effective potential morphology for electromagnetic perturbations with $s=1$, $\ell=6$ and $M=1$. The grey shaded region indicates the $\{\alpha,l_{0}\}$ parameter space where the potential exhibits a double-peak structure, a prerequisite for echoes. The blank region corresponds to a typical single-peak potential. The black curve shows the constraint $l_{0}=2M\alpha/(e^{2}+\alpha)$ where the hairy black hole’s horizon coincides with the seed Schwarzschild radius. The horizontal and vertical red dashed lines mark the potential varying along typical parameter paths analyzed in (b) and (c), respectively. (b): Typical potential profiles along the horizontal path from (a) at a fixed $\alpha=7$. This path demonstrates the transition from the Outer-Single-Peak regime (left end) to the Inner-Single-Peak regime (right end) as $\alpha$ increases. (c): Typical potential profiles along the vertical path from (a) at a fixed $l_{0}/M=0.94$. This path demonstrates the transition from the Inner-Single-Peak regime (bottom end) to the Outer-Single-Peak regime (top end) as $l_{0}$ increases.
@@
To further explore the phenomenological consequences of this transition, we must select specific paths through this parameter space for our subsequent QNM analysis. To this end, we choose two representative trajectories, one horizontal with fixed $l_{0}/M=0.94$ and one vertical with fixed $\alpha=7$, that traverse the borders between the single-peak and double-peak regions, as indicated by the red dashed lines in Fig. 1. Furthermore, as we see from Fig. 1 and Fig. 1 that the single peak regions can be further classified into Inner-Single-Peak and Outer-Single-Peak depending on which peak exists as the red dashed lines extend beyond the double-peak region. By analyzing the QNM spectra along these paths, we aim to uncover the distinctive phenomenological signatures associated with the emergence of the potential well, providing a direct link between the spacetime geometry and its observable ringdown properties.
@@ III Photon sphere and echo QNM families
Upon performing a temporal Fourier decomposition on the master scalar, $\Psi_{s,\ell,m}(t,r)=e^{-\mathrm{i}\omega t}\Psi_{s,\ell,m}(r)$, the master equation (3) transforms into a one-dimensional Schrödinger-like equation for the complex frequency $\omega$, namely
@@
$\Psi_{s,\ell,m}(r_{\star}\rightarrow-\infty)\sim e^{-\mathrm{i}\omega r_{\star}},\quad\text{and}\quad\Psi_{s,\ell,m}(r_{\star}\rightarrow\infty)\sim e^{\mathrm{i}\omega r_{\star}}.$ (9)
@@ III.2 Analysis of the QNM spectrum and method comparison
Using the methods listed above, we calculate the QNM spectra for selected parameter sets with fixed $\alpha=7$, showing in Tab. 1, which provides a detailed look at the QNM spectra for three representative cases and highlights the agreement between different computational methods. More importantly, it reveals the structural changes in the spectra as the effective potential transiting between single-peak and double-peak configurations.
@@
In contrast, for the Outer-Single-Peak case ($l_{0}/M=1.040$), the WKB fundamental mode matches the pseudospectral fundamental mode $(n=0)$. This scenario represents the familiar behavior expected from black holes with a typical potential with a single peak, where the spectrum is dominated by a single family of modes associated with the main potential barrier.
@@ III.3 Classification and evolution of QNM families
An interesting phenomenon occurs during the transition from the double peaks to the Inner-Single-Peak region. Unlike the abrupt disappearance of echo modes when transitioning to the Outer-Single-Peak region, the echo family extends smoothly into the Inner-Single-Peak domain. By a quick glance on Fig. 1, we argue that this is because the disappearance of the outer peak leaves behind a prominent “shoulder” in the effective potential. This shoulder creates a region where the potential decays slowly towards spatial infinity, leading to adiffraction trapping effect for waves. This underlying physical mechanism can be intuitively understand when one focus on the spacial configuration of a single mode, as it can be explicitly seen in Fig. 3, where we present the fundamental eigenfunction ($\Psi_{s,\ell,m}=\Psi_{1,6,0}$, blue curves) , computed by the direction integration method, against their related effective potentials ($V_{s}=V_{1}$, yellow curves) for each case.

### v2

# Quasinormal mode families: classification and competition

@@ Abstract
The perturbation spectra of black holes beyond standard vacuum black hole solutions within general relativity (GR) may exhibit complex structures with long-lived modes. This usually generates echo-like modulations on the ringdown signal, which typically originate from modified boundary conditions associated with exotic compact objects. Recent studies also reveal that they can instead be arised from the multi-peaked structure of the perturbation potential. However, while some case-by-case studies have been carried out, a framework for understanding the internal structure of such spectra, the physical nature of different mode families, and their dynamical excitation remains to be fully systematized. In this paper, we address this issue by proposing a potential methodology that combines frequency-domain classification with time-domain analysis, using a hairy Schwarzschild black hole that admits a double-peak perturbative potential as a theoretical platform Our analysis of the quasinormal mode (QNM) spectrum identifies two distinct families of modes: the photon sphere (PS) family, arising from delocalized scattering resonances, and the echo family, corresponding to highly localized quasi-bound states. We then develop a windowed energy analysis framework in the time domain, which discloses a dynamic competition for dominance between these families. In particular, our results explicitly show that this competition is sensitive to the properties of the initial perturbation source, and that higher-overtone echo modes can dominate in the observed signal, which are in contrast to the standard PS mode in GR. This study establishes the dynamic evolution of this energy competition as a new observational signature for probing new physics and further motivate a supplemental framework for analyzing long-lived ringdown signals.
@@ I Introduction
The dawn of the gravitational-wave (GW) astronomy era, inaugurated by the landmark observations of the Laser Interferometer Gravitational-Wave Observatory (LIGO) collaborations LIGOScientific:2016aoc , has opened an unprecedented window into the strong field regime of gravity. Future space-based detectors such as Laser Interferometer Space Antenna (LISA) Finn:2000sy , Taiji Hu:2017mde , and TianQin TianQin:2015yph and more promise to expand this window even further. One important scenario in this field is black hole spectroscopy Berti:2025hly ; Berti:2005ys , the idea that by analyzing the “ringdown” signal of a post-merger black hole, one can probe its intrinsic properties analogous to how atomic spectroscopy reveals the structure of atoms. Within the standard paradigm, the ringdown waveform is modeled as a superposition of QNMs Vishveshwara:1970zz ; Kokkotas:1999bd ; Berti:2009kk ; Konoplya:2011qq . These modes, characterized by their complex frequencies, are considered carriers of information about the final black hole’s mass and spin, providing a powerful framework to test the foundational no-hair theorem of GR Berti:2005ys ; Dreyer:2003bv ; Gossan:2011ha . Recent applications of black hole spectroscopy have been extended beyond the standard paradigm to incorporate nonlinearities Gleiser:1995gx ; Cheung:2023vki ; Redondo-Yuste:2023ipg ; Zhu:2024dyl , overtone contributions Giesler:2019uxc ; Bhagwat:2019dtm ; Baibhav:2023clw , and environmental effects Barausse:2014tra ; Cheung:2021bol ; Biswas:2023ofz . Furthermore, the intrinsic structure of the spectrum has been shown to exhibit rich features, including spectral instabilities Nollert:1996rf ; Jaramillo:2020tuu ; Cao:2024oud ; Cai:2025irl and resonant behaviors such as avoided crossings and exceptional points Onozawa:1996ux ; Cook:2014cta ; Motohashi:2024fwt ; Lo:2025njp ; Yang:2025dbn ; Cao:2025afs ; PanossoMacedo:2025xnf . Moreover, the scope of ringdown analysis has broadened to probe fundamental physics, such as spacetime symmetries Ghosh:2024het , non-gravitational modes arising from extra degrees of freedom Crescimbeni:2024sam , and so on.
@@
Within this expanding landscape, recent theoretical explorations have shown that the emergence of new QNM families is a widespread phenomenon in a variety of physical scenarios. These new families can be categorized by their physical origins. Firstly, modifications to spacetime’s asymptotic or topological structure can introduce new QNM families. A prime example is found in Cardoso:2017soq , where, in addition to the familiar photon sphere (PS) family tied to the black hole’s light ring, a distinct and long-lived de Sitter family arises from the presence of the cosmological horizon. Exotic compact objects like wormholes Cardoso:2016rao ; Bueno:2017hyj ; Qian:2025lau , or even specific braneworld models Tan:2024qij , can feature effective potential wells that trap perturbations, giving rise to long-lived echo families Cardoso:2019rvt . Additionally, the echo modes were also found and analyzed by imposing a modified boundary condition at the horizon Abedi:2020ujo ; Cardoso:2017cqb ; Cardoso:2016oxy or a non-smooth effective potential Liu:2021aqh ; Daghigh:2025wcw ; Xie:2025jbr . Secondly, the coupling between gravitational perturbations and other matter fields invariably leads to mixed spectra. For examples, the gravito-electromagnetic coupling in Kerr-Newman spacetimes splits the QNM spectra into PS and near-horizon families Dias:2015wqa ; and the gravitational and scalar modes in various scalar-tensor theories become intertwined Guo:2022rms ; Blazquez-Salcedo:2016enn ; Lin:2025lsz . The coupling within Einstein-Maxwell-scalar theories can also generate long-lived modes and associated echoes, a notable feature that has been shown to persist even at the fully nonlinear level Sang:2022hng ; Melis:2024kfr . Lastly, the intrinsic properties of the perturbing fields themselves, such as the mass of a Proca field, can generate quasi-bound states that form their own spectral families Rosa:2011my ; Lei:2023wlt . All these studies suggest that complex spacetime geometries or matter contents could leave their imprints directly on the global structure of the QNM spectra.
@@
Our methodology consists of a two-step workflow. The first step is a physical classification of modes in the frequency domain, which is made by WKB approximation, and we verify the reliability of the results of the QNM spectra with a set of methods. The second step is a dynamical analysis in time domain. To this end, we compute the complex amplitude of the master scalar Berti:2007dg ; Nollert:1998ys and develop a windowed energy analysis framework. By calculating the energy fraction of each mode as a function of parameters related with a sliding observation window, we construct a dynamic portrait of how different families compete for dominance and reveal the sensitivity of this competition to the initial perturbation source. It is noted that our proposed scheme can be applied for exploring QNM families in a broader studies associated with multi-peaked potentials. Here our analysis in Schwarzschild hairy black hole, somehow as an illustrative example, shall elaborate its application.
@@ II Perturbation equation and echo parameter space
$\text{with}\quad A(r)=B(r)=1-\frac{2M}{r}+\alpha\,\mathrm{e}^{-\frac{r}{M-l_{0}/2}}\equiv f(r)\quad\text{and}\quad C(r)=r^{2}\,.$ (2)
Here, $M$ represents the asymptotic mass of the hairy black hole, which is related with the seed Schwarzschild mass $\mathsf{M_{s}}$, via $M=\mathsf{M_{s}}+l_{0}/2$. The parameters $\{\alpha,l_{0}\}$ originate from the GD sector, where the positive $\alpha$ quantifies the deformation strength and $l_{0}$ corresponds to the charge associated with primary hair. More details can be seen from Appendix A. Note that $l_{0}\leq 2M$ should be satisfied to guarantee the asymptotic flatness of the hairy black hole Ovalle:2020kpd This black hole was found to admit an additional PS in certain parameter regimes Guo:2022ghl , comparing against the single PS in Schwarzschild black hole. These multiple PSs bring in richer and distinctive optical appearances Meng:2023htc ; Ribeiro:2025ohn and interferometric signatures Wang:2025hzu . The presence of additional PS has also found to closely associated with the echo signals of the probe scalar field Cavalcanti:2022cga ; Yang:2024rms . Moreover, the QNM and greybody factor for various perturbations of this black hole has been studied in Yang:2022ifo , in which the authors focus on the case with single peak in the perturbative potentials.
@@
where the field strength tensor $F^{\mu\nu}$ is defined in terms of the vector potential $A_{\mu}$ as $F=\mathrm{d}A$. In a spherically symmetric background, the vector potential can be decomposed into a Regge-Wheeler-Zerilli (RWZ) form Ruffini:1972pw ; Chandrasekhar:1985kt
@@
For polar modes with even parity $(-1)^{\ell}$, the master scalar can be taken as Chandrasekhar:1985kt
@@
We move on to investigate the morphology of the effective potentials. A crucial aspect of the hairy black hole solution (1) is the rich phenomenology that arises from its full parameter space. While the analysis in Cavalcanti:2022cga ; Avalos:2023jeh focused on a specific subspace by imposing the constraint $l_{0}=2M\alpha/(\mathrm{e}^{2}+\alpha)$, with $\mathrm{e}$ the base of the natural logarithm, which adjusts the horizon radius of the hairy black hole horizon radius to that of Schwarzschild black hole (e.g. $r_{h}(M,\alpha,l_{0})=2\mathsf{M}_{s}$). Here we shall explore the broader implications by treating the hairy charge $l_{0}$ and the deformation intensity $\alpha$} as independent variables. The interplay between these two parameters can induce a non-trivial structure in the spacetime geometry. Specifically, for certain regions in the $\{\alpha,l_{0}\}$ parameter space, the metric function $f(r)$ develops two inflection points outside the horizon. This non-monotonicity of the gravitational potential is directly imprinted onto the perturbative dynamics, causing the effective potential $V_{s}(r)$ to exhibit three local extrema.
@@
Figure 1: (a): Parameter spaces of the effective potential morphology for electromagnetic perturbations with $s=1$, $\ell=6$ and $M=1$. The grey shaded region indicates the $\{\alpha,l_{0}\}$ parameter space where the potential exhibits a double-peak structure, a prerequisite for echoes. The blank region corresponds to a typical single-peak potential. The black curve shows the constraint $l_{0}=2M\alpha/(\mathrm{e}^{2}+\alpha)$ where the hairy black hole’s horizon coincides with the seed Schwarzschild radius. The horizontal and vertical red dashed lines mark the potential varying along typical parameter paths analyzed in (b) and (c), respectively. (b): Typical potential profiles along the horizontal path from (a) at a fixed $l_{0}/M=0.94$. This path demonstrates the transition from the Outer-Single-Peak regime (left end) to the Inner-Single-Peak regime (right end) as $\alpha$ increases. (c): Typical potential profiles along the vertical path from (a) at a fixed $\alpha=7$. This path demonstrates the transition from the Inner-Single-Peak regime (bottom end) to the Outer-Single-Peak regime (top end) as $l_{0}$ increases. Note that in (b) and (c), the red squares mark an outer single peak which survives from the outer barrier of the double-peak phase as the parameter changes, while the green triangles mark an inner singe peak surviving from the inner barrier.
@@
We then scan the parameter $\{\alpha,l_{0}\}$ and map the morphology of the electromagnetic potential with fixed $s=1$, $\ell=6$ and $M=1$ (same settings for later studies) in Fig. 1, where the gray shaded region highlights the parameter combinations that give rise to the double-peak potential, in contrast to the blank region corresponding to the typical single-peak case. To further explore the phenomenological consequences of this transition, we must select specific paths through this parameter space for our subsequent QNM analysis. To this end, we choose two representative parameter trajectories that traverse the borders between the single-peak and double-peak regions, as indicated by the red dashed lines in Fig. 1: one horizontal with fixed $l_{0}/M=0.94$ and one vertical with fixed $\alpha=7$. The behavior of effective potential is shown in Figs. 1 and 1, respectively. In Fig. 1 with fixed $l_{0}/M=0.94$, we see that as $\alpha$ deceases, the inner peak marked by green triangle disappears while the outer peak marked by red Square survives, indicating the entering of Outer-Single-Peak regime. Inversely, when increasing $\alpha$, the outer peak disappears while the inner peak survives, indicating the entering of Inner-Single-Peak regime. Therefore, along the horizontal parameter trajectories with fixed $l_{0}/M=0.94$, the effective potential is in the Outer-Single-Peak regime, double-peak regime and Inner-Single-Peak regime as $\alpha$ increases. Similar phenomena is obvious for the vertical parameter trajectories with fixed $\alpha=7$, as shown in Fig. 1. Then by analyzing the QNM spectra along these paths, we aim to uncover the distinctive phenomenological signatures associated with the emergence of the potential well, providing a direct link between the spacetime geometry and its observable ringdown properties. As we will see soon, this classification of effective potentials based on the number of peaks, especially further tracing the outer and inner peak, is important for disclosing the global structure of the QNM spectrum.
@@ III Photon sphere and echo QNM families
Upon performing a temporal Fourier decomposition on the master scalar, $\Psi_{s,\ell,m}(t,r)=\mathrm{e}^{-\mathrm{i}\omega t}\Psi_{s,\ell,m}(r)$, the master equation (3) transforms into a one-dimensional Schrödinger-like equation for the complex frequency $\omega$, namely
@@
$\Psi_{s,\ell,m}(r_{\star}\rightarrow-\infty)\sim\mathrm{e}^{-\mathrm{i}\omega r_{\star}},\quad\text{and}\quad\Psi_{s,\ell,m}(r_{\star}\rightarrow\infty)\sim\mathrm{e}^{\mathrm{i}\omega r_{\star}}.$ (9)
@@ III.2 Analysis of the QNM spectrum and method comparison
Using the methods listed above, we calculate the QNM spectra for selected parameter sets with fixed $\alpha=7$, showing in Tab. 1, which provides a detailed look at the QNM spectra for three representative cases and highlights the agreement between different computational methods. More importantly, it reveals the structural changes in the spectra as the effective potential transiting between single-peak and double-peak configurations. Specifically, the discrepancy between WKB predictions and pseudospectral results manifests differently depending on the potential’s profile. The pseudospectral method, validated by the matrix pencil method, proves essential for identifying the full spectrum where the WKB approximation serves as a preliminary mode classifier. We discuss these specific manifestations below:
@@
  2. 2.
    For the Outer-Single-Peak case ($l_{0}/M=1.040$), the WKB fundamental mode matches the pseudospectral fundamental mode $(n=0)$. This scenario represents the familiar behavior expected from black holes with a typical potential with a single peak, where the spectrum is dominated by a single family of modes associated with the main potential barrier.
  3. 3.
@@ III.3 Classification and evolution of QNM families
An interesting phenomenon occurs during the transition from the double peaks to the Inner-Single-Peak region. Unlike the abrupt disappearance of echo modes when transitioning to the Outer-Single-Peak region, the echo family extends smoothly into the Inner-Single-Peak domain. By a quick glance on Fig. 1, we argue that this is because the disappearance of the outer peak leaves behind a prominent “shoulder” in the effective potential. This shoulder creates a region where the potential decays slowly towards spatial infinity, leading to a diffraction trapping effect for waves. Noted that the existence of such modes has also been independently found in a recent work Wang:2025mxe in which the authors identified them as the off-peak modes. This underlying physical mechanism can be intuitively understand when one focus on the spacial configuration of a single mode, as it can be explicitly seen in Fig. 3, where we present the fundamental eigenfunction ($\Psi_{s,\ell,m}=\Psi_{1,6,0}$, blue curves) , computed by the direction integration method, against their related effective potentials ($V_{s}=V_{1}$, yellow curves) for each case.